OpenStudy (anonymous):

15. A recursive function is shown below: f(1) = 2 and f(n) = f(n – 1) – 5; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? (4 points) A. 2, 3, 8, 13, 18, ... B. 2, 7, 12, 17, 22, ... C. 2, –3, –8, –13, –18, ... D. 2, –7, –12, –17, –22, ...

OpenStudy (anonymous):

If \(f(1)=2\), then \(f(2)=f(2-1)-5=f(1)-5=2-5=-7\). Only one of the options has 2, -7, ...

OpenStudy (anonymous):

Oh okay; Thanks very much

OpenStudy (anonymous):

yw